OpenStudy (anonymous):

Evaluate the function for the given value of x. f(x) = -6x2 - 4x - 4 , x = - 2 A. f( - 2) = -36 B. f( - 2) = -17 C. f( - 2) = -20 D. f( - 2) = -28 I am really stuck on this question and my teacher isn't online too help me. Obviously it isn't D and I don't think it's A I need help!?

OpenStudy (anonymous):

Plug your value in for x

OpenStudy (anonymous):

@kelseymariah

OpenStudy (anonymous):

it will be -20 when the x value is negative 2

OpenStudy (anonymous):

Hmm give me a sec

OpenStudy (anonymous):

That's not what I'm getting

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

you're subtracting a negative by a negative so it will be -16 plus -4 which is -20 :)

OpenStudy (anonymous):

It's -20, yes. Would you like a step by step?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

I got -20 but I always like to check

OpenStudy (anonymous):

After you plug in your values, you get +6x^2-4x-4 = -6x(-2)^2- -2*4-4

OpenStudy (anonymous):

-6x^2**

OpenStudy (anonymous):

Evaluate (-2)^2 = 4 ; +6*4- -2*4-4

OpenStudy (anonymous):

-6*4 = -24- -2*4-4

OpenStudy (anonymous):

-24+8-4

OpenStudy (anonymous):

8-(24+4) = 8-28 = -20

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yep!